# Latent Diffusion Inversion Requires Understanding the Latent Space

### v1

@@ 4.1 Setup
For CIFAR‑10 and CelebA [liu2015deep], we trained a $\beta$-VAE [higgins2017beta] from scratch on the member set. Then we use the latent representations of member set to train a latent space diffusion model unconditionally while VAE encoder-decoder is frozen. From each training split we subsample $n$ images and partition them equally into a member set and a held‑out set. ImageNet-1K [russakovsky2015imagenet] is trained at two different points. (1) We directly employ the publicly released autoencoder model stabilityai/sd-vae-ft-mse [stabilityai_sdvae_ft_mse] without further fine-tuning. This model, developed by Stability AI, serves as the standard latent autoencoder used in the Stable Diffusion v1 framework. (2) The latent space diffusion model is also trained from scratch. We train it conditonally by using the architecture components from diffuser, which we guarantee have a similar architecture to vanilla LDM [rombach2022high]. The KL-regularization ($\beta$) for three datasets are 1e21e-2, 1e31e-3, 1e61e-6. With higher resolution, one expects to reduce the $\beta$ to ensure the same scale of regularization across datasets. and The architecture details are released in Appendix G.
@@
For Pokémon[huggingface_pokemon], COCO2017‑Val [lin2014microsoft], and Flickr30k [young2014image], we fine‑tuned Stable Diffusion v1‑4111https://huggingface.co/CompVis/stable-diffusion-v1-4 on a randomly selected subset of each training split, reserving an equally sized subset as the held‑out set. None of these datasets are in the original pre‑training corpus.
Stable Diffusion v1‑4 was fine‑tuned for 15k, 100k, and 200k steps on the member splits of Pokémon, COCO‑2017‑Val, and Flickr30k, respectively, with a fixed learning rate of $1\times 10^{-5}$ (AdamW). Additionally, we adopted the default data augmentation (Random-Crop and Random-Flip) while training.
4.2 Experiments on local distortion vs. memorization
To compute the local distortion at each latent point, we employ our matrix-free randomized SVD procedure (Appendix 1). This algorithm evaluates Jacobian–vector products using jvp/vjp primitives in PyTorch, thereby avoiding explicit construction of the full decoder Jacobian. When automatic differentiation is unavailable, we find that finite-difference approximations provide a sufficiently accurate substitute for jvp. Unless otherwise stated, we use fixed randomized SVD hyperparameters across all experiments: target rank $k=20$, oversampling parameter $p=30$, and $q=2$ power iterations, which balance computational efficiency and numerical stability. For measuring memorization, we primarily adopt SimA [rao2025score] due to its simplicity and effectiveness.
@@
Across all datasets, we observe a consistent and pronounced ordering: higher-distortion quartiles yield substantially higher attack AUCs, indicating that regions of large decoder distortion exhibit stronger memorization. The magnitude of this effect varies by dataset. For CIFAR-10, the distortion gap between quartiles is large (Fig. 2), producing correspondingly large separations in attack performance; in contrast, CelebA exhibits nearly uniform distortion throughout its latent space, resulting in only minor differences in AUC across quartiles. We leave MNIST high/low distortion samples in the appendix. Interestingly, the lines on ImageNet-1K experiments cross at one point. This weird crossing is dataset-dependent. Given that the earlier time MIA is not trustworthy [rao2025score], it still shows a consistent phenomenon after passing the early time. This strong correlation between decoder’s local distortion and memorization indicates that the choice of encoder–decoder architecture plays a central role in shaping the learned score field and, consequently, the memorization behavior of latent diffusion models.
@@ 4.3 Experiments on per-dimension memorization
To justify our work, we selectively choose four state-of-the-art threat models: SimA [rao2025score], SecMI [duan2023diffusion], PIA [kong2023efficient], and Loss (also refered to “Naive” method) [matsumoto2023membership] as our benchmark methods acknowledging that list is not exhaustive. For each method, we followed the hyperparameter suggestion in their original paper. Notably, $l_{4}\text{-norm}$, $l_{2}\text{-norm}$, $l_{4}\text{-norm}$ and $l_{2}\text{-norm}$ were used for SimA, SecMI, PIA and Loss as suggested.
The complete results are reported in Tables 1 and 2. We report the best performance over probe time $t=[0:300]$. Throughout all experiments, we apply a dilution procedure, which removes the top-$k\%$ least influential latent dimensions from the attack vector prior to computing the norm. Although we do not conduct a search for the optimal value of $k$, we expect that such an optimal choice may vary across datasets and individual samples; exploring this dependency is an interesting direction for future work. For simplicity and consistency, we thus adopt $k=40$ as a global setting. As a control, we also include a baseline that randomly discards $40\%$ of dimensions when applying SimA (first row in each table), which typically reduces attack performance. In contrast, removing the least-memorizing dimensions based on our proposed per-dimension metric consistently improves all three evaluation measures across all methods and datasets. And from the “min $\delta$”, we found that Loss method usually achives the least improvement compared to others. These results indicate that latent-space diffusion models exhibit dimension-wise non-uniform memorization, with certain latent coordinates contributing disproportionately to overfitting and membership leakage.
@@ 4.4 Connection to high-frequency components
lian2025unveiling reported that reconstructing images from denoised noise vectors and suppressing high-frequency components made members and non-members more separable. This approach performed well for diffusion models trained directly in the data domain, but in our experiments it reduced performance for more complex latent-space diffusion models (e.g., Stable Diffusion); see Section E in the appendix. We hypothesize that if the latent encoder were perfectly linear, then manipulating Fourier components would have behaved as expected in latent space. However, the non-linear VAE encoder used in latent diffusion models does not simply map images to Fourier modes; instead, it performs denoising on a sub-manifold of the data domain for which the spectral analog is generally inaccessible (the spectrum of the Laplace-Beltrami operator).
To test this hypothesis, we computed the correlation between the local distortion of each image and the summed energy of its low-frequency and high-frequency spectra, for each of the three datasets CIFAR-10, CelebA, and ImageNet-1K. To select low frequencies, we summed the center elements of the Fourier transformed data; for CIFAR-10 ($32^{2}$), CelebA ($64^{2}$) and ImageNet-1K ($256^{2}$) the central radii summed were 5, 10, 40, which are approximately $2.5\%$ of the spectra for each. The results, summarized in Table 3, show that at a high level the distortion reflects the intuition of lian2025unveiling, but that this analog is imperfect, especially for more structured datasets such as CelebA, where all data have aligned structure (spatially registered faces, in the case of CelebA).
@@ 5 Discussion
Limitations and Caveats:
We have tested only a subset of attack methods and a subset of possible datasets. Conditionalization, dataset size, and dataset structure clearly influence diffusion model behavior and thereby inversion dynamics. We can clearly design datasets where the data domain is similar to these latent spaces, and we can design trivial “VAEs” with no curvature. Our results are likely relevant for similar image data, which represent the mainstay of diffusion model applications, but they may not generalize to disparate domains (e.g., discrete diffusion on text or tabular data).
Crossover strata:
The ImageNet subfigure of Figure 3 (at far right) show that the distortion strata lines cross over each other around timestep 70. This seems to inform upon interactions between the noise schedule and the length-scale of the latent space; clearly for faster noise schedules this cross-over point would occur earlier. Interactions between noise schedules and decoder geometry may be important for attack parameter tuning and overall understanding of overfit phenomena.
@@ 6 Conclusion
[…]
@@ A Approximation of Local Distortion
B Plots of attack score versus local distortion
This is repetitive content. We keep it here to maintain the indexing consistency between the main text and the appendix. It will be deleted in future versions.
@@ D Per–Dimension Influence via Hutchinson Estimator
E Latent diffusion attack by smoothing high-frequency components
In this section, we apply the method proposed by lian2025unveiling to latent diffusion models (LDMs), specifically Stable Diffusion. Although the method was originally developed for data-domain diffusion models, we extend it to the latent domain using the following formulation.
@@ E Latent diffusion attack by smoothing high-frequency components
where $\epsilon_{\theta}(\cdot)$ indicates the predicted noise. According to Appendix B of lian2025unveiling. We derive
@@
where $\epsilon_{\theta}(z_{0},0)$ represents the noise prediction for $z_{0}$. According to Appendix B of lian2025unveiling, it can be converted to
@@
A high-frequncy filter is defined the same as li2024unveiling,
@@
Table 5: Attack performance of four baseline methods used to evaluate memorization of Stable Diffusion. Best results in bold.
@@
F More image examples on different local distortion
More examples of images in the high (low) distortion region are provided in Fig.  4.
@@ F More image examples on different local distortion
G Architecture details of all LDM instances

### v3

@@ 4.1 Setup
For CIFAR‑10 and CelebA [21], we trained a $\beta$-VAE [15] from scratch on the member set. Then we use the latent representations of member set to train a latent space diffusion model unconditionally while VAE encoder-decoder is frozen. From each training split we subsample $n$ images and partition them equally into a member set and a held‑out set. ImageNet-1K [29] is trained at two different points: (1) We directly employ the publicly released autoencoder model stabilityai/sd-vae-ft-mse [1] without further fine-tuning. This model, developed by Stability AI, serves as the standard latent autoencoder used in the Stable Diffusion v1 framework. (2) The latent space diffusion model is also trained from scratch. We train it conditonally by using the architecture components from diffuser, which we guarantee have a similar architecture to vanilla LDM [28]. The KL-regularization ($\beta$) for three datasets are $10^{-2}$, $10^{-3}$, $10^{-6}$. With higher resolution, one expects to reduce the $\beta$ to ensure the same scale of regularization across datasets. and The architecture details are released in Appendix H. We also implement the proposed method on LDMs with VQ-GAN [10] as the encoder (See details in Appendix E).
@@
For Pokémon[7], COCO2017‑Val [20], and Flickr30k [36], we fine‑tuned Stable Diffusion v1‑4111https://huggingface.co/CompVis/stable-diffusion-v1-4 on a randomly selected subset of each training split, reserving an equally sized subset as the held‑out set. None of these datasets are in the original pre‑training corpus.
Stable Diffusion v1‑4 was fine‑tuned for 15k, 100k, and 200k steps on the member splits of Pokémon, COCO‑2017‑Val, and Flickr30k, respectively, with a fixed learning rate of $10^{-5}$ (AdamW). Additionally, we adopted the default data augmentation (Random-Crop and Random-Flip) while training.
4.2 Experiments on distortion and memorization
To compute the local distortion at each latent point, we employ our matrix-free randomized SVD procedure (Appendix 1). This algorithm evaluates Jacobian-vector products using jvp/vjp primitives in PyTorch, thereby avoiding explicit construction of the full decoder Jacobian. When automatic differentiation is unavailable, we find that finite-difference approximations provide a sufficiently accurate substitute for jvp. Unless otherwise stated, we use fixed randomized SVD hyperparameters across all experiments: target rank $k=20$, oversampling parameter $p=30$, and $q=2$ power iterations, which balance computational efficiency and numerical stability. For measuring memorization, we primarily adopt SimA [27] due to its simplicity and effectiveness.
@@ 4.2 Experiments on distortion and memorization
Across all datasets, we observe a consistent and pronounced ordering: higher-distortion quartiles yield substantially higher attack AUCs, indicating that regions of large decoder distortion exhibit stronger memorization. The magnitude of this effect varies by dataset. For CIFAR-10, the distortion gap between quartiles is large (Fig. 2), producing correspondingly large separations in attack performance; in contrast, CelebA exhibits nearly uniform distortion throughout its latent space, resulting in only minor differences in AUC across quartiles. We leave MNIST high/low distortion samples in the appendix. Interestingly, the lines on ImageNet-1K experiments cross at one point. This weird crossing is dataset-dependent. Given that the earlier time MIA is not trustworthy [27], it still shows a consistent phenomenon after passing the early time. This strong correlation between decoder’s local distortion and memorization indicates that the choice of encoder–decoder architecture plays a central role in shaping the learned score field and, consequently, the memorization behavior of latent diffusion models.
@@ 4.3 Experiments on per-dimension memorization
To justify our work, we selectively choose four state-of-the-art threat models: SimA [27], SecMI [9], PIA [17], and Loss (also refered to “Naive” method) [23] as our benchmark methods acknowledging that list is not exhaustive. For each method, we followed the hyperparameter suggestion in their original paper. Notably, $l_{4}\text{-norm}$, $l_{2}\text{-norm}$, $l_{4}\text{-norm}$ and $l_{2}\text{-norm}$ were used for SimA, SecMI, PIA and Loss as suggested.
The complete results are reported in Tables 1 and 2. We report the best performance over probe time $t=[0:300]$. Throughout all experiments, we apply a dilution procedure, which removes the top-$k\%$ least influential latent dimensions from the attack vector prior to computing the norm. Although we do not conduct a search for the optimal value of $k$, we expect that such an optimal choice may vary across datasets and individual samples; exploring this dependency is an interesting direction for future work. For simplicity and consistency, we thus adopt $k=40$ as a global setting. As a control, we also include a baseline that randomly discards $40\%$ of dimensions when applying SimA (first row in each table), which typically reduces attack performance (A table of various filtering ratios $k$ is provided in the Appendix B). In contrast, removing the least-memorizing dimensions based on our proposed per-dimension metric consistently improves all three evaluation measures across all methods and datasets. And from the “min $\delta$”, we found that Loss method usually achives the least improvement compared to others. These results indicate that latent-space diffusion models exhibit dimension-wise non-uniform memorization, with certain latent coordinates contributing disproportionately to overfitting and membership leakage.
@@ 4.4 Connection to high-frequency components
Lian et al. [19] reported that reconstructing images from denoised noise vectors and suppressing high-frequency components made members and non-members more separable. This approach performed well for diffusion models trained directly in the data domain, but in our experiments it reduced performance for more complex latent-space diffusion models (e.g., Stable Diffusion); see Section F in the appendix. We hypothesize that if the latent encoder were perfectly linear, then manipulating Fourier components would have behaved as expected in latent space. However, the non-linear VAE encoder used in latent diffusion models does not simply map images to Fourier modes; instead, it performs denoising on a sub-manifold of the data domain for which the spectral analog is generally inaccessible (the spectrum of the Laplace-Beltrami operator).
To test this hypothesis, we computed the correlation between the local distortion of each image and the summed energy of its low-frequency and high-frequency spectra, for each of the three datasets CIFAR-10, CelebA, and ImageNet-1K. To select low frequencies, we summed the center elements of the Fourier transformed data; for CIFAR-10 ($32^{2}$), CelebA ($64^{2}$) and ImageNet-1K ($256^{2}$) the central radii summed were 5, 10, 40, which are approximately $2.5\%$ of the spectra for each. The results, summarized in Table 3, show that at a high level the distortion reflects the intuition of Lian et al. [19], but that this analog is imperfect, especially for more structured datasets such as CelebA, where all data have aligned structure (spatially registered faces, in the case of CelebA).
@@ 5 Discussion
Limitations and Caveats: We have tested only a subset of attack methods and a subset of possible datasets. Conditionalization, dataset size, and dataset structure clearly influence diffusion model behavior and thereby inversion dynamics. We can clearly design datasets where the data domain is similar to these latent spaces, and we can design trivial “VAEs” with no curvature. Our results are likely relevant for similar image data, which represent the mainstay of diffusion model applications, but they may not generalize to disparate domains (e.g., discrete diffusion on text or tabular data).
Crossover strata: The ImageNet subfigure of Figure 3 (at far right) show that the distortion strata lines cross over each other around timestep 70. This seems to inform upon interactions between the noise schedule and the length-scale of the latent space; clearly for faster noise schedules this cross-over point would occur earlier. Interactions between noise schedules and decoder geometry may be important for attack parameter tuning and overall understanding of overfit phenomena.
@@ 6 Conclusion
7 Acknowledgements
This work was supported in part by NSF 2321684 and the ARPA-H ALISS project.
@@ A Approximation of Local Distortion
B Variation of Filtering Ratio $k$
We evaluated masking ratios $k\in[10\%,90\%]$ on ImageNet-1K(subset) of SimA [27] attack statistics. The results are summarized on Table 4
Table 4: AUC performance of SimA attack under different masking ratios $k$ on ImageNet-1K (subset). For comparison, the last column presents a reverse-masking setting where the top 90% most-memorizing dimensions are removed. Retaining only the bottom 10% renders the attacks almost ineffective.
Masking Ratio $k$ 0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 0.9 (Rev)
AUC ($\uparrow$) 69.70 70.77 71.68 72.32 72.64 72.54 71.95 70.65 68.91 65.99 54.67
TPR@1%FPR ($\uparrow$) 3.96 4.68 5.64 7.06 7.79 8.09 8.62 9.04 8.46 6.78 2.21
@@ D Per–Dimension Influence via Hutchinson Estimator
E Experiments on LDMs with VQ-GAN as the encoder
For experiments on MNIST, CIFAR-10 and CelebA, we used a vanilla $\beta$-VAE. For experiments on ImageNet, Pokémon, COCO‑2017‑Val, and Flickr30k, we employed Stability AI’s autoencoders trained with Adversarial (GAN) and Perceptual (LPIPS) losses. We also test our masking strategy on VQ-GAN [10](discrete, large geometry change). The original LDM implementation used the bottleneck layer $\hat{z}$ before quantization; since we need the decoder side as well, we approximate $\frac{\partial D(\hat{z})}{\partial\hat{z}}\approx\frac{\partial D(z_{q})}{\partial z_{q}}$ after quantization. We trained a VQ-GAN from scratch for CIFAR-10 and MNIST and used a pre-trained one [^2] https://ommer-lab.com/files/latent-diffusion/vq-f8.zip for ImageNet-1K. Training details, such as the number of training steps, are kept consistent with the other experiments. Table 6 shows that our method consistently looks for the most-memorizing dimension, which corresponds to the per-dimensional distortion.
Table 6: LDMs with VQGAN as the encoder. Performance (AUC %) comparison on different datasets. Numbers in bold indicate improvement.
Method MNIST CIFAR-10 ImageNet-1K
SimA (After Filter) 83.69 (+2.7) 86.27 (+3.2) 67.38 (+3.85)
F Latent diffusion attack by smoothing high-frequency components
In this section, we apply the method proposed by Lian et al. [19] to latent diffusion models (LDMs), specifically Stable Diffusion. Although the method was originally developed for data-domain diffusion models, we extend it to the latent domain using the following formulation.
@@ F Latent diffusion attack by smoothing high-frequency components
where $\epsilon_{\theta}(\cdot)$ indicates the predicted noise. According to Appendix B of Lian et al. [19]. We derive
@@
where $\epsilon_{\theta}(z_{0},0)$ represents the noise prediction for $z_{0}$. According to Appendix B of Lian et al. [19], it can be converted to
@@
A high-frequncy filter is defined the same as Li et al. [18],
@@
Table 7: Attack performance of four baseline methods used to evaluate memorization of Stable Diffusion.
@@
G More image examples on different local distortion
More examples of images in the high (low) distortion region are provided in Fig.  4. In a standard VAE architecture, the latent space is a bottleneck, meaning $k<d$. If $x=D(z)$ is purely deterministic, the generated data does not span the full $\mathbb{R}^{d}$ space. Instead, it lies strictly on a $k$-dimensional manifold embedded within the $d$-dimensional ambient space.
Because of this, $p(x)$ is a degenerate distribution in $\mathbb{R}^{d}$ (it has zero volume) and the standard change of variables formula fails. However, if we evaluate the density strictly on the manifold with respect to the $k$-dimensional Hausdorff measure, we must use the area formula. The metric tensor induced on the latent space by the mapping $D$ is $J_{D}(z)^{T}J_{D}(z)$.The density on the manifold is [4, 12]:
$p(x)=p(z)\left(\det\left(J_{D}(z)^{T}J_{D}(z)\right)\right)^{-1/2}$
Taking the logarithm gives:
$\log p(x)=\log p(z)-\frac{1}{2}\log\underbrace{\det\left(J_{D}(z)^{T}J_{D}(z)\right)}_{\text{Distortion(z)}}$ (18)
According to [26, 24, 30], the log-likelihood estimation in a generative model is highly biased toward the input complexity, where the model assigns more density to complex samples. Based on Eq. 18, this insight explains the positive correlation between visual complexity and the distortion of a sample in Fig. 2 and Fig. 4. Hence, the distribution shifts significantly from the data space to the latent space.
@@ G More image examples on different local distortion
H Architecture details of all LDM instances